# When and How Does CLIP Enable Domain and Compositional Generalization?

### v1

@@ Abstract
The remarkable generalization performance of contrastive vision-language models like CLIP is often attributed to the diversity of their training distributions. However, key questions remain unanswered: Can CLIP generalize to an entirely unseen domain when trained on a diverse mixture of domains (domain generalization)? Can it generalize to unseen classes within partially seen domains (compositional generalization)? What factors affect such generalization? To answer these questions, we trained CLIP models on systematically constructed training distributions with controlled domain diversity and object class exposure. Our experiments show that domain diversity is essential for both domain and compositional generalization, yet compositional generalization can be surprisingly weaker than domain generalization when the training distribution contains a suboptimal subset of the test domain. Through data-centric and mechanistic analyses, we find that successful generalization requires learning of shared representations already in intermediate layers and shared circuitry.
@@ 1 Introduction
Recent work has converged on the conclusion that CLIP’s diverse training distribution is the primary factor driving its unprecedented generalization performance. For example, Fang et al. (2022) found that other factors such as language supervision, training data size, or the contrastive loss play only a minor role, while Nguyen et al. (2022) showed that data quality is more important than quantity. More recently, Mayilvahanan et al. (2024b) demonstrated that CLIP’s “generalization performance […] drops to levels similar to what has been observed for ImageNet-trained models” (Mayilvahanan et al., 2024b, p. 10) by limiting the diversity of (visual) domains [^1] We refer to a domain as a group of images sharing a common style, such as natural images, sketches, paintings, etc. to a minimum, i.e., by removing all non-natural samples. This shows that the mixture of various (non-natural) domains plays an important role for CLIP’s generalization, yet the underlying mechanisms remain unexplored. This brings us to our core research question:
@@
In particular, we investigate under which circumstances CLIP can learn the object class invariances across the training domains with the aim to generalize to entirely unseen domains–a fundamental question about its domain generalization capability (Blanchard et al., 2011; Muandet et al., 2013; Gulrajani & Lopez-Paz, 2021). We also study questions about CLIP’s compositional generalization (Hupkes et al., 2020; Wiedemer et al., 2023), which is believed to be an important factor of its generalization performance (Mayilvahanan et al., 2024a; Udandarao et al., 2024) and a long-standing challenge of machine learning research. Adapting Szabó’s (2012) classical example, we ask pictorially: Can CLIP, trained on natural images of cats and dogs along with sketches of cats, generalize to sketches of dogs?
@@
  • Compositional generalization is challenging: Surprisingly, including a domain in the training data does not always improve generalization to unseen classes within that domain. However, ensuring sufficient class diversity within the test domain–ideally with no overlap with the queried classes in evaluation–along with high domain diversity, can significantly reduce the aforementioned performance gap.
  • Generalization requires sufficient feature and circuit sharing: When CLIP generalizes well compositionally, it shares more embeddings and intermediate features between different domains. However, CLIP sometimes fails to generalize to certain domains. We provide a twofold explanation: (1) the inputs from these domains lack shared features and, as a result, (2) the model has limited shared intermediate features and circuitry between domains, constraining its ability to generalize effectively. We support this hypothesis through representational similarity analysis and introduce a related concept for circuits [^2] Interconnected internal model mechanisms/components for performing a specific computation or task (Olah et al., 2020).: mechanistic similarity, which measures similarity between circuits.
This work presents a comprehensive study for systematically investigating CLIP’s domain generalization and compositional generalization capabilities. This enables us to unveil the capabilities and limitations of CLIP. Our code will be published together with the conference paper.
@@ 2 Related Work
CLIP’s remarkable OOD generalization has sparked research in identifying the factors driving it. Fang et al. (2022) showed that CLIP’s training distribution–rather than its dataset size, language supervision, or contrastive loss–is the primary factor of its OOD generalization. Similarly, Nguyen et al. (2022) found that dataset quality outweighs quantity. However, the precise characteristics of the training distribution contributing to CLIP’s generlization performance remained unclear.
While high train-test similarity was initially believed to be a key factor, Mayilvahanan et al. (2024a) found its impact to be smaller than expected. Instead, other factors, such as the class distribution, were shown to have a more important role (Wen et al., 2024). Moreover, caption richness has been found to enhance CLIP’s robustness (Xue et al., 2024; Wen et al., 2024), and CLIP’s loss fosters the learning of disentangled representations, facilitating the generalization to unseen attribute-object combinations (Abbasi et al., 2024). At the same time, other work revealed that CLIP exhibits behaviors resembling those of supervised classifiers. For example, CLIP fails to generalize when all non-natural images are removed from its training data (Mayilvahanan et al., 2024b), and CLIP can be vulnerable to spurious correlations (Wang et al., 2024). While these works have significantly advanced our understanding of CLIP’s OOD generalization, key questions remain. For example, “Can CLIP generalize to an entirely unseen domain?” (domain generalization), “Can CLIP generalize to unseen class-domain combinations?” (compositional generalization), and which factors contribute to such generalization?
@@ 3 Problem Setup
Figure 1: Training data setups and CLIP’s performance in these setups. A: We systematically varied the domain mixture and object class exposure of CLIP’s training data across four scenarios while controlling for other factors like dataset size and model choice. Specifically, we trained CLIP models on (1) mostly natural images (“Natural-only”) to obtain a lower performance bound, (2) a diverse set of domains excluding the test domain (“Leave-out-domain”) to assess its domain generalization, (3) natural images with a subset of test domain classes (“CG low-diversity”) to evaluate its compositional generalization, and (4) a combination of diverse domains plus a subset of test domain classes (“CG high-diversity”) for a more diverse compositional generalization setting. B: CLIP trained on diverse domains demonstrates stronger OOD generalization compared to training on only natural images or fewer domains. Remarkably, CLIP can perform as well or better even without exposure to some part (i.e., object classes) of the test domain during training.
Recent work highlighted that CLIP is trained on a substantial amount of non-natural images (Mayilvahanan et al., 2024b, Table 2), but the role of these non-natural images in enabling CLIP’s OOD generalization remains unclear. In this work, we aim to address this question by systematically analyzing the effect of different domain mixtures in CLIP’s training data (see Figure 1), allowing us to study CLIP’s ability to generalize to entirely unseen domains (domain generalization) and to novel combinations of known domains and classes (compositional generalization).
@@
Let $D_{0}$ denote the base domain mostly consisting of natural images with image-text pairs $(I^{0}_{i},T^{0}_{i})$ (red in Figure 1). Further, we consider $m$ non-natural domains $D_{r}$ for $r\in\{1,\dots,m\}$ with image-text pairs $(I^{r}_{i},T^{r}_{i})$ (blue, green, orange). Lastly, we consider the object classes $C=\{c_{1},\dots,c_{n}\}$ for the images $I$ which we divide into two disjoint subsets $C_{1}=\{c_{1},\dots,c_{k}\}$ (squares) and $C_{2}=\{c_{k+1},\dots,c_{n}\}$ (circles). We denote the subset of $D_{r}$ that contains only classes from $C^{\prime}$ as $D_{r}^{C^{\prime}}$.
@@
  • Natural-only (lower bound): We train only on our base image domain $D_{0}$ (e.g., ImageNet-Captions (Fang et al., 2022), see Section 4 for further details), consisting (almost) exclusively of natural images of all classes $C$. This condition serves as our lower bound and mirrors the training distributions considered in Fang et al. (2022); Mayilvahanan et al. (2024b).
  • Leave-out-domain (domain generalization): We train on a diversity of domains $\bigcup_{j\neq i}D_{j}\cup D_{0}$ but hold out the test domain $D_{i}$, following the classical domain generalization learning setup (Blanchard et al., 2011; Muandet et al., 2013; Gulrajani & Lopez-Paz, 2021). Note we could also interpret this as extrapolation in certain cases, i.e., does CLIP generalize to data outside the domain coverage seen during training?
@@
A model compositionally generalizes if it can accurately classify any new combinations of seen factors (here, classes and domains) not seen together during training.
@@
  • CG low-diversity: We train on the base domain $D_{0}$ and a subset of classes $C_{1}$ of the test domain $D_{i}^{C_{1}}$: $D_{0}\cup D_{i}^{C_{1}}$.
  • CG high-diversity: We train on the base domain $D_{0}$, a subset of classes $C_{1}$ of the test domain $D_{i}^{C_{1}}$, and a diverse set of other domains $\mathbf{D}_{j\neq i}:=\bigcup_{j\neq i}D_{j}$ containing all classes $C$: $D_{0}\cup D_{i}^{C_{1}}\cup\mathbf{D}_{j\neq i}$.
@@
Our test data for all training data setups consists of the same novel combinations of the classes $C_{2}$ of the test domain $D_{i}$: $D_{i}^{C_{2}}$. By keeping the test set fixed throughout all conditions, we ensure comparability across these setups.
@@ 5 When Does CLIP Exhibit Domain and Compositional Generalization?
In this section, we trained CLIP models on our systematically constructed training data setups (refer to Section A.6 for training details), as described in the previous sections and illustrated in Figure 1, to investigate when CLIP can achieve domain and compositional generalization. Figure 2 summarizes the results for CLIP models with ResNet-50 vision encoder and trained on ImageNet-Captions as base dataset. We discuss the results and key findings below.
To ensure the validity of our results and findings, we validated them across several alternative choices: (1) base datasets (CC3M, CC12M), (2) vision encoders (ViT-S-32, Swin-T), and (3) contrastive loss choices (SigLIP (Zhai et al., 2023)). These additional results, which are consistent with those in Figure 2, are provided in Section B.1.
@@
Figure 2: High diversity domain mixtures exhibit improved effective robustness. Each point represents the average performance over three consecutive training epochs and three seeds, with higher opacity indicating later training epochs. High diversity domain mixtures, such as Leave-out-domain (red) and CG high diversity (purple), have consistently higher generalization performance than their low diversity counterparts. These gains are especially pronounced in the clipart and sketch domains. However, for the quickdraw domain, generalization fails even in the high diversity settings–a limitation we will further investigate in Section 6.2.
@@
Figure 3: Certain domains contribute more strongly to generalization performance than others. When domains are added successively from CG low-diversity (left) to CG high-diversity (right), some domains significantly improve generalization, while others provide only small gains or even slightly degrade performance.
Figure 3 shows that certain domains contribute more strongly than others, while other domains can even slightly decrease performance on the test domain. These observations suggest that while domain diversity is generally beneficial, the relationship between the added domain(s) and the test domain is a critical factor in facilitating generalization. {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 1: Domain diversity enhances domain generalization and compositional generalization. However, the overlap between domains matters, too.
@@
Table 1 shows that, even in high diversity settings with better generalization, a gap to the upper bound performance remains. We analyzed the number of test class samples required to close this gap in Section B.3. We found that the number of required samples tends to scale linearly.
Table 1: There is a performance gap when CLIP has seen close class samples of the test domain. With sufficient domain diversity, CLIP generalizes well to unseen classes $C_{2}$ of the clipart and sketch test domains $D_{i}$. However, even in these cases, there remains a gap compared to models also trained on domain-specific samples of the classes, i.e., $D^{C_{2}}_{i}$. Section B.3 provides the results of the other domains.
@@
While high diversity improves generalization, one may expect CLIP to generalize better to unseen classes within a test domain if it has seen some other classes of that domain during training. However, Figure 2 surprisingly reveals the opposite: CLIP models trained on a subset of classes from the test domain (CG low-/high-diversity) are often slightly outperformed by models that have not seen the test domain at all (Natural-only, Leave-out-domain). {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 2: Compositional generalization can be weaker than domain generalization. This finding is surprising, since the test domain is entirely unseen and domain generalization can require extrapolation. In contrast, compositional generalization is expected to perform better as it has partial exposure to that domain. However, our results suggest that this intuition may not always hold.
To better understand this, we investigated the role of the test domain’s chosen classes for training and the ones that are queried during evaluation. For example, CLIP may learn the shortcut that all sketches belong to the subset of seen classes, which becomes wrong for the unseen classes queried in evaluation. To test this hypothesis, we replaced DomainNet’s sketches of the classes $C_{1}$ in our previous CG settings with sketches of the classes $C^{\prime}_{1}$ that are not queried in evaluation. For this, we used ImageNet-Sketch (Wang et al., 2019) and excluded all ImageNet classes that overlap with the classes from DomainNet; refer to Section A.7 for details on the class overlap.
Table 2: Seeing a subset of classes of the test domain can worsen compositional generalization. Including a subset of sketches from DomainNet (CG low/high-diversity) slightly decreases performance on the unseen sketch classes $C_{2}$ compared to not seeing that domain at all (Leave-out-domain). However, adding sketches of classes that do not overlap with DomainNet’s classes, instead improves compositional generalization performance, suggesting that compositional generalization can be limited by (only) partial, suboptimal inclusion of the test domain.
@@
To investigate the influence of language supervision, we replicated our previous experiments with a supervised classifier. Specifically, we trained the classifier on the combined class distribution of ImageNet and DomainNet. Refer to Section A.7 for further (training) details.
Figure 4: Effective robustness of CLIP vs. supervised classifiers. Similar to CLIP, the robustness of supervised classifiers also increases with the domain diversity of the training data. However, CLIP consistently shows superior performance. Refer to Section B.4 for the results on the remaining domains.

### v2

@@ Abstract
The remarkable generalization performance of contrastive vision-language models like CLIP is often attributed to the diversity of their training distributions. However, key questions remain unanswered: Can CLIP generalize to an entirely unseen domain when trained on a diverse mixture of domains (domain generalization)? Can it generalize to unseen classes within partially seen domains (compositional generalization)? What factors affect such generalization? To answer these questions, we trained CLIP models on systematically constructed training distributions with controlled domain diversity and object class exposure. Our experiments show that domain diversity is essential for both domain and compositional generalization, yet compositional generalization can be surprisingly weaker than domain generalization when the training distribution contains a suboptimal subset of the test domain. Through data-centric and mechanistic analyses, we find that successful generalization requires the learning of sufficiently shared representations in intermediate layers and circuits.
@@ 1 Introduction
Recent work has converged on the conclusion that CLIP’s diverse training distribution is the primary factor driving its unprecedented generalization performance. For example, Fang et al. (2022) found that other factors such as language supervision, training data size, or the contrastive loss play only a minor role, while Nguyen et al. (2022) showed that data quality is more important than quantity. More recently, Mayilvahanan et al. (2025) demonstrated that CLIP’s “generalization performance […] drops to levels similar to what has been observed for ImageNet-trained models” (Mayilvahanan et al., 2025, p. 10) by limiting the diversity of (visual) domains [^1] We refer to a domain as a group of images sharing a common style, such as natural images, sketches, paintings, etc. to a minimum, i.e., by removing all non-natural samples. This shows that the mixture of various (non-natural) domains plays an important role for CLIP’s generalization, yet the underlying mechanisms remain unexplored. This brings us to our core research question:
@@
In particular, we investigate under which circumstances CLIP can learn the object class invariances across the training domains with the aim to generalize to entirely unseen domains—a fundamental question about its domain generalization capability (Blanchard et al., 2011; Muandet et al., 2013; Gulrajani & Lopez-Paz, 2021). We also study questions about CLIP’s compositional generalization (Hupkes et al., 2020; Wiedemer et al., 2023), which is believed to be an important factor of its generalization performance (Mayilvahanan et al., 2024; Udandarao et al., 2024) and a long-standing challenge of machine learning research. Adapting Szabó’s (2012) classical example, we ask pictorially: Can CLIP, trained on natural images of cats and dogs along with sketches of cats, generalize to sketches of dogs?
Figure 1: Training data setups and CLIP’s generalization performance under varying domain mixtures and class exposure. A: We systematically varied the domain mixture and object class exposure in CLIP’s training data across four controlled scenarios, while keeping other factors like dataset size and model choice constant. Specifically, we trained CLIP models on (1) Natural-only—mostly natural images (from classes $\blacksquare$, ) to establish a lower bound on performance; (2) Leave-out-domain—a diverse set of domains ($\blacksquare$, , $\blacksquare$, , $\blacksquare$, ) excluding the test domain to assess CLIP’s domain generalization; (3) CG low-diversity—natural images ($\blacksquare$, ) with some test domain classes ($\blacksquare$) to evaluate its compositional generalization with limited domain diversity; and (4) CG high-diversity—a diverse mix of domains ($\blacksquare$, , $\blacksquare$, , $\blacksquare$, ) plus some test domain classes ($\blacksquare$) to assess compositional generalization with higher domain diversity. B: CLIP models trained on diverse domains demonstrate stronger OOD generalization to held-out test domain classes (), compared to models trained only on natural images or fewer domains. Notably, CLIP can perform as well or better even without exposure to some test domain classes during training.
@@
  • Compositional generalization is challenging: Surprisingly, including a domain in the training data does not always improve generalization to unseen classes within that same domain. However, ensuring sufficient class diversity within the test domain—ideally with no overlap with the queried classes in evaluation—along with high domain diversity, can significantly reduce the aforementioned performance gap.
  • Generalization requires sufficient feature and circuit sharing: When CLIP generalizes well compositionally, it shares more embedding features between different domains. However, CLIP sometimes fails to generalize to certain domains. We provide a twofold explanation: (1) the inputs from these domains lack shared input features and, as a result, (2) the model has limited shared intermediate features and circuits [^2] Interconnected internal model mechanisms/components for performing a specific computation or task (Olah et al., 2020). between domains, constraining its ability to generalize effectively. We support this hypothesis through representational similarity analysis and introduce a related concept for circuits: mechanistic similarity, measuring similarity between circuits.
This work presents a systematic study of CLIP’s domain and compositional generalization, unveiling both its capabilities and its limitations. For reproducibility, our code is available at https://github.com/lmb-freiburg/understanding-clip-ood.
@@ 2 Related Work
Unlike earlier models from the ImageNet era, CLIP models have been shown to exhibit remarkable OOD generalization (Radford et al., 2021; Geirhos et al., 2021). Subsequent work has further enhanced this capability by introducing stylistic diversity or encouraging the learning of domain-invariant representation (Huang et al., 2023; Bose et al., 2024; Addepalli et al., 2024; Yu et al., 2024). These advances in OOD generalization have sparked research to uncover the underlying factors behind CLIP’s strong OOD generalization.
Fang et al. (2022) showed that CLIP’s diverse training distribution—rather than its dataset size, language supervision, or contrastive loss—is the primary factor of its OOD generalization. Similarly, Nguyen et al. (2022) found that dataset quality outweighs quantity. However, the precise characteristics of the training distribution contributing to CLIP’s generlization performance remained unclear.
While high train-test similarity was initially believed to be a key factor, Mayilvahanan et al. (2024) found its impact to be smaller than expected. Instead, other factors, such as the class distribution, were shown to play a more important role (Wen et al., 2024). Moreover, caption richness has been found to enhance CLIP’s robustness (Xue et al., 2024; Wen et al., 2024), and CLIP’s loss fosters the learning of disentangled representations, facilitating the generalization to unseen attribute-object combinations (Abbasi et al., 2024). At the same time, other work revealed that CLIP exhibits behaviors resembling those of supervised classifiers. For example, CLIP fails to generalize when all non-natural images are removed from its training data (Mayilvahanan et al., 2025), and CLIP can be vulnerable to spurious correlations (Wang et al., 2024). While these works have significantly advanced our understanding of CLIP’s OOD generalization, key questions remain. For example, “Can CLIP generalize to an entirely unseen domain?” (domain generalization), “Can CLIP generalize to unseen class-domain combinations?” (compositional generalization), and which factors contribute to such generalization?
@@ 3 Problem Setup
Recent work highlighted that CLIP is trained on a substantial amount of non-natural images (Mayilvahanan et al., 2025, Table 2), but the role of these non-natural images in enabling CLIP’s OOD generalization remains unclear. In this work, we aim to address this question by systematically analyzing the effect of different domain mixtures in CLIP’s training data (see Figure 1), allowing us to study CLIP’s ability to generalize to entirely unseen domains (domain generalization) and to novel combinations of known domains and classes (compositional generalization).
@@
Let $D_{0}$ denote the base domain mostly consisting of natural images with image-text pairs $(I^{0}_{i},T^{0}_{i})$ (red in Figure 1). Further, we consider $m$ non-natural domains $D_{r}$ for $r\in\{1,\dots,m\}$ with image-text pairs $(I^{r}_{i},T^{r}_{i})$ (blue, green, orange). Lastly, we consider the object classes $C=\{c_{1},\dots,c_{n}\}$ for the images $I$ which we divide into two disjoint subsets $C_{1}=\{c_{1},\dots,c_{k}\}$ (squares $\blacksquare$) and $C_{2}=\{c_{k+1},\dots,c_{n}\}$ (circles ⚫). We denote the subset of $D_{r}$ that contains only classes from $C^{\prime}$ as $D_{r}^{C^{\prime}}$.
@@
  • Natural-only (lower bound): We train only on our base image domain $D_{0}$ (e.g., ImageNet-Captions (Fang et al., 2022), see Section 4 for further details), consisting (almost) exclusively of natural images of all classes $C$ ($\blacksquare$, ). This condition serves as our lower bound and mirrors the training distributions considered in Fang et al. (2022); Mayilvahanan et al. (2025).
  • Leave-out-domain (domain generalization): We train on a variety of domains but hold out the test domain $D_{i}$: $\bigcup_{j\neq i}D_{j}\cup D_{0}$ ($\blacksquare$, , $\blacksquare$, , $\blacksquare$, ); following the classical domain generalization learning setup (Blanchard et al., 2011; Muandet et al., 2013; Gulrajani & Lopez-Paz, 2021). Note we could also interpret this as extrapolation in certain cases, i.e., does CLIP generalize to data outside the domain coverage seen during training?
@@
A model compositionally generalizes if it can accurately classify any new combinations of seen factors—here, classes and domains—not seen together during training.
@@
  • CG low-diversity: We train on the base domain $D_{0}$ and a subset of classes $C_{1}$ of the test domain $D_{i}$: $D_{0}\cup D_{i}^{C_{1}}$ ($\blacksquare$, , $\blacksquare$).
  • CG high-diversity: We train on the base domain $D_{0}$, a diverse set of other domains $\mathbf{D}_{j\neq i}:=\bigcup_{j\neq i}D_{j}$ containing all classes $C$, and a subset of classes $C_{1}$ of the test domain $D_{i}$: $D_{0}\cup\mathbf{D}_{j\neq i}\cup D_{i}^{C_{1}}$ ($\blacksquare$, , $\blacksquare$, , $\blacksquare$, , $\blacksquare$).
@@
Our test data for all training data setups consists of the same novel combinations of the classes $C_{2}$ of the test domain $D_{i}$: $D_{i}^{C_{2}}$ (). By keeping the test set fixed throughout all conditions, we ensure comparability across these setups.
@@ 5 When Does CLIP Exhibit Domain and Compositional Generalization?
In this section, we trained CLIP models on our systematically constructed training data setups, as described in the previous sections and illustrated in Figure 1 (refer to Section A.6 for training details), to investigate when CLIP achieves domain generalization and compositional generalization. Figure 2 summarizes the results for CLIP models with ResNet-50 vision encoder and trained on ImageNet-Captions as base dataset. We discuss the results and key findings below.
To ensure the validity of our results and findings, we validated them across several alternative choices: (1) base datasets (CC3M, CC12M), (2) vision encoders (ViT-S-32, Swin-T), (3) contrastive loss choices (SigLIP; Zhai et al., 2023), and (4) vision-language pre-training methods (BLIP; Li et al., 2022). These additional results, which are consistent with those in Figure 2, are provided in Section B.1.
@@
Figure 2: High diversity domain mixtures exhibit improved effective robustness. Each point represents the average balanced top-1 accuracy over three consecutive training epochs and three seeds, with higher opacity indicating later training epochs. High diversity domain mixtures (Leave-out-domain (red), CG high-diversity (purple)) have consistently higher generalization performance than their low diversity counterparts (Natural-only (blue), CG low-diversity (green)). These gains are especially pronounced in the Clipart and Sketch domains (leftmost and rightmost). However, for the Quickdraw domain (4th from left), generalization fails even in the high diversity settings—a limitation we will further investigate in Section 6.2. Other evaluation metrics (balanced top-5 accuracy and $F_{1}$ score) yield consistent results (see Figure 8 in Appendix B).
@@
(a) Clipart.
(b) Sketch.
Figure 3: Certain domains contribute more strongly to generalization performance than others. When domains are added successively from CG low-diversity to CG high-diversity (from left to right), some domains significantly improve generalization performance (balanced top-1 accuracy) on the unseen classes of the test domain, while others provide only small gains or even slightly degrade it.
Figure 3 shows that some domains contribute more strongly to generalization, while other domains can even slightly reduce generalization performance on the test domain. Intriguingly, the impact of a domain can change, depending on which other domains are already included. For example, in Figure 3(a), while Sketch and Quickdraw are the most beneficial domains to include initially for improving generalization performance on Clipart, once Sketch is added, it is more effective to first include the other domains (Painting, Infograph) before adding Quickdraw. We hypothesize that this is due to the visual similarity between Sketch and Quickdraw—both consist of black-and-white/gray images with object contours—which may lead the model to learn similar features. Once such features are learned from one domain, the benefit of adding the other diminishes. Further investigation of this observation is left for future work.
Overall, these observations suggest that while domain diversity is generally beneficial for generalization, the specific relationship between the added domain(s) and the test domain plays a critical role. {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 1: Domain diversity enhances domain generalization and compositional generalization. However, the relationship between domains also matters.
@@
Table 1 shows that, even in high diversity settings with better generalization, a gap to this upper bound performance remains. We analyzed the number of test class samples required to close this gap in Section B.3. We found that the number of required samples tends to scale linearly.
Table 1: There is a performance gap when CLIP has seen class-specific samples of the test domain. With sufficient domain diversity, CLIP generalizes well to unseen the classes $C_{2}$ of the Clipart and Sketch test domains $D_{i}$. However, even in these cases, there remains a gap in balanced top-1 accuracy compared to models also trained on domain-specific samples of the classes, i.e., $D^{C_{2}}_{i}$. Section B.3 provides the results of the other domains.
@@
While high diversity improves generalization, one may intuitively expect CLIP to generalize better to unseen classes within a test domain if it has seen some other classes of that domain during training. However, Figure 2 surprisingly reveals the opposite: CLIP models trained on a subset of classes from the test domain (CG low-/high-diversity) are often slightly outperformed by models that have not seen the test domain at all (Natural-only, Leave-out-domain). {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 2: Compositional generalization can be weaker than domain generalization. This finding is surprising, since the test domain is entirely unseen and domain generalization can require extrapolation. In contrast, compositional generalization is expected to perform better as it has partial exposure to that domain. However, our results suggest that this intuition may not always hold.
To better understand this, we investigated the role of the test domain’s chosen classes for training and the ones that are queried during evaluation. For example, CLIP may learn the shortcut that all sketches belong to the subset of the seen classes, which becomes wrong for the unseen classes queried in evaluation. To test this hypothesis, we replaced DomainNet’s sketches from the classes $C_{1}$ in our previous CG settings with sketches from a set of classes $C_{3}$ which is disjoint from $C_{1}\cup C_{2}$ and not queried in evaluation. For this, we used ImageNet-Sketch (Wang et al., 2019) and excluded all ImageNet classes that overlap with the classes from DomainNet; refer to Section A.3 for details on the class overlap.
Table 2: Seeing a subset of classes of the test domain can worsen compositional generalization. Including a subset of sketches from DomainNet (CG low-/high-diversity) slightly decreases balanced top-1 accuracy on the unseen sketch classes $C_{2}$ compared to not seeing that domain at all (Leave-out-domain). However, adding sketches of classes that do not overlap with DomainNet’s classes, instead improves compositional generalization performance, suggesting that compositional generalization can be limited by (only) partial, suboptimal inclusion of the test domain.
@@
To investigate the influence of language supervision, we replicated our previous experiments with a supervised ResNet50 classifier. Specifically, we trained the classifier on the combined class distribution of ImageNet and DomainNet. Refer to Section A.7 for further training details.
Figure 4: Effective robustness of CLIP vs. supervised classifiers. Similar to CLIP, the robustness of supervised classifiers also increases with the domain diversity of the training data. However, CLIP consistently shows superior balanced top-1 accuracy. Refer to Section B.4 for the results on the remaining domains.